explain how to add and subtract rational expressions with different denominators. Provide examples to support your explanation.
Answer by Kiersten(3)

1/4+2/8
you have to find common denom. which number will go both into 4 and 8 .4 will . 4 will go into 4 once and 8 twice.

1/4+2/8
2/8+2/8=4/8=1/2